OpenStudy (anonymous):

How do you simplify this 3(2+√ 3)

OpenStudy (anonymous):

expand 3x2 + 3 x sqrt3 = \[6 + 3\sqrt{3}\]

OpenStudy (anonymous):

ask your teacher why \[6+3\sqrt{3}\] is simpler than \[3(2+\sqrt{3})\]